In three years time Joanne will be twice as old as her daughter. If Joanne is 45 years old now how old is her daughter?
Answer by Cromlix(4381)

In three years time Joanne will be twice as old as her daughter.
J + 3 = 2(D + 3)
If Joanne is 45 years old now
45 + 3 = 2(D + 3)
48 = 2D + 6
2D = 48 - 6
2D = 42
D = 21
Daughter is 21 years old.
Hope this helps.
;-)